OpenStudy (anonymous):

expressing this product as a sum cos(5η) cos(3η)

OpenStudy (anonymous):

would it be (1/2)[cos(2η) + cos(8η)]

OpenStudy (anonymous):

these type of problems, you can always google for

OpenStudy (anonymous):

\[\cos(A-B)=\cos A \cos B+\sin A \sin B ...(1)\] \[\cos \left( A+B \right)=\cos A \cos B-\sin A \sin B ...(2)\] adding (1) and (2)

OpenStudy (anonymous):

\[2\cos A \cos B=\cos \left( A-B \right)+\cos \left( A+B \right)\]

OpenStudy (anonymous):

yea so it should be (1/2)[cos(2η) + cos(8η)]

OpenStudy (anonymous):

correct.